OpenStudy (anonymous):

HELP! Giving out medals what is the solution for x^2 +6x=-5

OpenStudy (anonymous):

x^2+6x+5=0 x^2+5x+x+5 x(x+5)+1(x+5)=0 (x+1)(x+5)=0

OpenStudy (anonymous):

x = -1 or -5

OpenStudy (anonymous):

Thanks Guys!! c: